# Uncovering LLM-Generated Code: A Zero-Shot Synthetic Code Detector via Code Rewriting

### v1

@@ 5.2. Effectiveness of Detector LLM
We observed that using GPT-3.5-turbo as the detector LLM for rewriting achieves optimal performance in detecting code generated by StarChat, GPT-3.5-turbo, and GPT-4, except for CodeLlama. Optimal performance in detecting code generated by CodeLlama is attained when using CodeLlama as the LLM detector.The primary reason for this is that among the four models (CodeLlama, StarChat, GPT-3.5-turbo and GPT-4), StarChat is fine-tuned on an instruction dataset distilled from GPT-3.5-turbo and GPT-4. Therefore, these three models may share a closer distribution. In contrast, CodeLlama is instruction-tuned on a dataset constructed by self-instructing Llama-2 (Touvron et al., 2023). We suspect that the suboptimal performance in detecting CodeLlama-generated code using GPT-3.5-turbo is due to the differences in the instruction datasets’ distributions.
Regarding the optimal use of GPT-3.5-turbo over StarChat for detecting StarChat-generated code, our observation of the rewritten code generated by both GPT-3.5-turbo and StarChat revealed that StarChat tends to oversimplify the rewritten code in complex problems, while GPT-3.5-turbo provides more accurate rewritten code.
@@ 5.3. Comparison to Supervised Detector
In Table 3, the first two rows represent two supervised detectors, GPTZero and OpenAI-Detector. We compare them with the zero-shot detectors and find that GPTZero and OpenAI-Detector fail to provide useful detection for code content, performing similarly to random guessing. Despite being trained on millions of labeled samples, these supervised detectors may only capture a small fraction of code content. This observation suggests that supervised detectors are prone to overfitting the training distribution and struggle to generalize to new domains without adaptive tuning. Furthermore, the proposed zero-shot detection methods demonstrate promising generalization capability when evaluated on different code distributions and outperform supervised detectors.
@@ 5.4. Ablation Study
We conducted an ablation experiment to analyze the contributions of two primary components in our design, i.e., Code Rewriting and Similarity Measurement. We considered two ablation setting: first, replacing the Code Rewriting with in-fill perturbation following DetectGPT (Mitchell et al., 2023) while retaining the Similarity Measurement (Sim); second, retaining Code Rewriting (CR) and replacing the similarity measurement between original and rewritten code with token-wise score difference (i.e, LogProb, LogRank, Rank and Entropy). The results are presented in Table 4.
@@ 6. Model Analysis
We conduct comprehensive experiments to assess the generalization of the proposed methods under diverse settings, including the impact of generation prompts, decoding strategy, programming languages, code correctness, code length, and revised scenarios. In addition, we also provide a sensitivity test for the choice of $m$ and the similarity models to understand our methods’ effectiveness better. For all experiments in this section, we use GPT-3.5-turbo as the generation tool.
@@ 6.1. Choice of Similarity Model
In the method section, we exploit self-supervised contrastive learning (Gao et al., 2021) to train a better code similarity model based on GraphCodeBERT (GCB-SimCSE). To investigate the impact of the similarity model, we experiment with three other variants of code similarity models: UniXcoder (Guo et al., 2022) trained by SimCSE (Unix-SimCSE), the original GraphCodeBERT model with average pooling (GCB-avg) and OpenAI’s text embedding service (Text-ada-002) (Greene et al., 2022). The results on the APPS benchmark (generated by GPT-3.5-turbo) are listed in Table 5. The UniXcoder trained with self-supervised contrastive learning (Unix-SimCSE) attains the highest performance, yet using the other three alternatives still surpasses the previous zero-shot detectors. The experimental results indicate that our method is not dependent on one specific similarity model. Moreover, as UniXcoder is an enhanced version of GraphCodeBERT, the superior performance of Unix-SimCSE over our original GCB-SimCSE implies that the detection performance can be further improved by utilizing a more strong and robust similarity model.
@@
Table 5. Test with different similarity models. Text-ada-002 is OpenAI’s text embedding service for text/code search or clustering. We use CodeLlama, StarChat and GPT-3.5-turbo as Detector LLMs.
@@ 6.2. Impact of Generation Prompts
It is well known that the output of large language model can be significantly affected by prompts. Consequently, the prompts used to generate synthetic code have a substantial impact on the generated output. To evaluate the impact of generation prompts, we altered the prompts used to generate synthetic code in our benchmark by removing the chain of thought prompts and using direct prompting, i.e., asking the model to generate a solution directly. The difference between these two prompts are illustrated in Figure 4 and the results under two different prompts are presented in Figure 5.
@@ 6.3. Impact of Decoding Strategy
Adjusting the temperature parameter in large language models balances output diversity and accuracy, with lower temperatures yielding more deterministic and consistent results, while higher temperatures produce more varied and creative outputs. We conducted an experiment on APPS benchmark where we varied the generator temperature across the range of $[0.2,0.4,0.8]$ while keeping the rewriting temperature constant at $0.8$. The results are shown in Figure 6. This temperature range was chosen based on the findings from Codex paper (Chen et al., 2021), which identified $T=0.2$ and $T=0.8$ were optimal for pass@1 and pass@100 rates, respectively. The results in Figure 6 demonstrate that our method exhibits superior consistency across different temperatures.
@@ 6.4. Generalizability to Different Programming Languages
To verify the generalizability of our method to different programming languages, we construct an additional C++ benchmark using the Code Contest dataset (Li et al., 2022b). We generate synthetic code using the same four-generation tools. The detection performance on Code Contest C++ is presented in Table 6. Our proposed method still achieves notable improvement compared to other zero-shot baselines on C++ benchmarks. Additionally, we observe that synthetic C++ code is easier to detect for both our method and other baselines when compared to Python, suggesting a more significant distribution gap between synthetic and human-written C++ code.
@@ 6.5. Detecting Revised Synthetic Code
In real-world scenarios, humans may revise synthetic code before using it. Naturally, this raises the question of whether zero-shot synthetic code detectors remain effective when a human user makes minor revisions to the generated code. We consider the most straightforward code modification: identifier renaming. This does not affect the functionality of the original code and even does not require the user to understand the code’s logic. To mimic identifier renaming, we extract all the identifier names and randomly replacing $10\%$, $20\%$, and $50\%$ fraction with ”$var\_i$”. Then we evaluate all the zero-shot detectors on the revised synthetic code to observe how performance changes with increasing fractions of replaced identifiers. The results are plotted in Figure 7.
@@
From Figure 7, it’s evident that the performance of all detectors declines as the replacement fraction increase, degrading to random guess when $50\%$ of the identifiers are replaced. However, our method consistently outperforms all other detectors across all replacement fraction levels. This phenomenon can be attributed to the behavior of GPT-3.5-turbo and StarChat when rewriting the revised synthetic code. They tend to restore ”$var\_i$” to variable names that contain code semantics, resulting in a lower similarity between the rewritten code and the revised synthetic code compared to the similarity between the rewritten code and the original synthetic code. We consider detection in this or more complex adversarial scenarios as a focus for future work.
@@ 6.6. Impact of Code Correctness
We consider that the distribution of correct code solutions likely differ from incorrect ones since the correct model-generated code is much closer to human-written code and is more difficult to detect. To explore the impact of code correctness, we separately present the detection AUROC for both correct and incorrect codes on MBPP benchmark when using CodeLlama as Detector LLM in Figure 8.
@@
From the results, we observe that both with our method and other baselines, detecting correct code is more challenging than detecting incorrect code. Nonetheless, our method consistently outperforms other baselines when it comes to correct code.
@@ 6.7. Impact of Code Length
We also investigate how the length of code snippet impacts the detection performance of the zero-shot synthetic code detectors. We divide the samples in the APPS benchmark into five groups according to the code string length. Each group has an equal number of synthetic and human-written code snippets. The results are plotted in Figure 9.
@@
In general, the performance of all detectors improves as the code length falls within the 0-1000 character range. The $\log p(x)$, LogRank, and Entropy methods achieve optimal detection performance on code snippets between 600-1000, while exhibiting inferior performance on code lengths smaller than 200. DetectGPT’s performance surpasses ours when the code exceeds 1000 characters. However, DetectGPT is inferior to the other baselines for code length less than 600, the range where majority ($70\%$) of code snippets lie. Our method displays strong performance across all ranges and is more robust to changes in code length when compared to the other baselines.
@@ 6.8. Choice of $m$
In our primary experiment, we set the number of rewriting $m$ in the range of $[2,4,8]$ due to limited computational resources. However, increasing $m$ can reduce randomness in code sampling and enhance the accuracy of expectation estimation. To investigate this, we conduct experiments by setting the maximum value of $m$ to 32 on the APPS and MBPP benchmarks (generated by GPT-3.5-turbo) and plotted the detection AUROC against changes of $m$ in Figure 10. The result indicate that the detection performance increases with $m$ with slight fluctuation and saturate around 20 - 32 rewrites.
@@ 7. Conclusion
In this paper, we identified the gap when applying synthetic content detectors designed for general texts to code content, and proposed a novel zero-shot synthetic code detector via code rewriting and similarity measurement. We utilize the similarity between rewritten and original code as an indicator for detecting synthetic code. Our proposed method addresses the unique challenges posed by the massive ”low-entropy” tokens in code content. Through exhaustive experiments, we uncovered several findings regarding synthetic code detection. The proposed method demonstrates superior performance and robustness in detecting synthetic code compared to other zero-shot baselines.

### v2

@@ 5.2. Effectiveness of Detector LLM
We observe that using GPT-3.5-turbo as the detector LLM for rewriting achieves optimal performance in detecting code generated by StarChat, GPT-3.5-turbo, and GPT-4, except for CodeLlama. Optimal performance in detecting code generated by CodeLlama is attained when using CodeLlama self as the LLM detector. The primary reason for this is that among the four models (CodeLlama, StarChat, GPT-3.5-turbo, and GPT-4), StarChat is fine-tuned on an instruction dataset distilled from GPT-3.5-turbo and GPT-4. Therefore, these three models may share a closer distribution. In contrast, CodeLlama is instruction-tuned on a dataset constructed by self-instructing Llama-2 (Touvron et al., 2023). We suspect that the suboptimal performance in detecting CodeLlama-generated code using GPT-3.5-turbo is due to the differences in the instruction datasets’ distributions.
Regarding the optimal use of GPT-3.5-turbo over StarChat for detecting StarChat-generated code, our observation of the rewritten code generated by both GPT-3.5-turbo and StarChat revealed that StarChat tends to oversimplify the rewritten code in complex problems, while GPT-3.5-turbo provides more accurate and complete rewritten code.
@@ 5.3. Comparison to Supervised Detector
In Table 3, the first two rows represent two supervised detectors, GPTZero and OpenAI-Detector. We compare them with the zero-shot detectors and find that GPTZero and OpenAI-Detector fail to provide useful detection for code content, performing similarly to random guessing. Despite being trained on millions of labeled samples, these supervised detectors may only capture a small fraction of code content. This observation suggests that supervised detectors are prone to overfitting the training distribution and struggle to generalize to new domains without adaptive tuning. Furthermore, our proposed zero-shot detection methods demonstrate promising generalization capability when evaluated on different code distributions and outperform supervised detectors.
@@ 5.4. Ablation Study
We conducted an ablation experiment to analyze the contributions of two primary components in our design, i.e., Code Rewriting and Similarity Measurement. We considered two ablation settings: first, replacing the Code Rewriting with in-fill perturbation following DetectGPT (Mitchell et al., 2023) while retaining the Similarity Measurement (Sim); second, retaining Code Rewriting (CR) and replacing the Similarity Measurement between original and rewritten code with token-wise score difference (i.e, LogProb, LogRank, Rank and Entropy). The results are presented in Table 4.
@@ 6. Model Analysis
We conduct comprehensive experiments to evaluate our proposed methods across diverse settings and scenarios. These experiments encompass a range of factors, including the choices of the Similarity Model, the impact of generation prompts, decoding strategy, generalizability to different programming languages, detection of revised synthetic code, consideration of code correctness, and assessment of code length impact. Furthermore, we perform a sensitivity analysis for the choice of $m$ to understand our methods’ effectiveness better. For all experiments in this section, we use GPT-3.5-turbo as the generation tool.
@@ 6.1. Choice of Similarity Model
In the Methodology section, we exploit self-supervised contrastive learning (Gao et al., 2021) to train a better code similarity model based on GraphCodeBERT (GCB-SimCSE). To investigate the impact of the Similarity Model, we experiment with three other variants of code similarity models: UniXcoder (Guo et al., 2022) trained by SimCSE (Unix-SimCSE), the original GraphCodeBERT model with average pooling (GCB-avg) and OpenAI’s text embedding service (Text-ada-002) (Greene et al., 2022). The results on the APPS benchmark (generated by GPT-3.5-turbo) are listed in Table 5. The Unix-SimCSE attains the highest performance, yet using the other three alternatives still surpasses the previous zero-shot detectors. The experimental results indicate that our method is not dependent on one specific similarity model. Moreover, as UniXcoder is an enhanced version of GraphCodeBERT, the superior performance of Unix-SimCSE over our original GCB-SimCSE implies that the detection performance can be further improved by utilizing a more strong and robust similarity model.
@@
Table 5. Test with different similarity models. We use CodeLlama, StarChat and GPT-3.5-turbo as Detector LLMs.
@@ 6.2. Impact of Generation Prompts
It is widely acknowledged that LLM outputs can be notably influenced by prompts. Therefore, the prompts used to generate synthetic code also have a substantial impact on the generated output. To assess this impact, we modified the prompts used in our benchmark for synthetic code generation. Specifically, we compared the effectiveness of two prompting approaches: chain of thought prompts and direct prompts. The latter directly asks the model to generate a solution without preliminary analysis. The difference between these two prompts is illustrated in Figure 4 and the results under two different prompts are presented in Figure 5.
@@ 6.3. Impact of Decoding Strategy
Adjusting the temperature parameter in LLMs balances output diversity and accuracy, with lower temperatures yielding more deterministic and consistent results, while higher temperatures produce more varied and creative outputs. To investigate this, we conducted an experiment on the APPS benchmark where we varied the generator temperature across the range of $[0.2,0.4,0.8]$ while keeping the rewriting temperature constant at $0.8$. The results are shown in Figure 6. This temperature range is chosen based on the findings from Codex paper (Chen et al., 2021), which identified $T=0.2$ and $T=0.8$ were optimal for pass@1 and pass@100 rates, respectively. The results in Figure 6 demonstrate that our method exhibits superior consistency across different temperatures.
@@ 6.4. Generalizability to Different Programming Languages
To assess the generalizability of our method to different programming languages, we construct an additional C++ benchmark using the Code Contest dataset (Li et al., 2022b). We generate synthetic code using the same four-generation tools. The detection performance on Code Contest C++ is presented in Table 6. Our method still achieves notable improvement compared to other zero-shot baselines on C++ benchmarks. Moreover, we observe that synthetic C++ code is easier to detect for both our method and other baselines when compared to Python, suggesting a more significant distribution gap between synthetic and human-written C++ code.
@@ 6.5. Detecting Revised Synthetic Code
In real-world scenarios, humans may revise synthetic code before using it. Naturally, this raises the question of whether zero-shot synthetic code detectors remain effective when a human user makes minor revisions to the generated code. We consider the most straightforward code modification: identifier renaming. This does not affect the functionality of the original code and even does not require the user to understand the code’s logic. To mimic identifier renaming, we extract all the identifier names and randomly replace $10\%$, $20\%$, and $50\%$ fraction with ”$var\_i$”. Then we evaluate all the zero-shot detectors on the revised synthetic code to observe how performance changes with increasing fractions of replaced identifiers. The results are plotted in Figure 7.
@@
From Figure 7, it’s evident that the performance of all detectors declines as the replacement fraction increases, degrading to random guess when $50\%$ of the identifiers are replaced. However, our method consistently outperforms all other detectors across all replacement fraction levels. This phenomenon can be attributed to the behavior of GPT-3.5-turbo and StarChat when rewriting the revised synthetic code. They tend to restore ”$var\_i$” to variable names that contain code semantics, resulting in a lower similarity between the rewritten code and the revised synthetic code compared to the similarity between the rewritten code and the original synthetic code. We consider detection in this or more complex adversarial scenarios as a focus for future work.
@@ 6.6. Impact of Code Correctness
We consider that the distribution of correct code solutions likely differs from incorrect ones since the correct synthetic code is much closer to human-written code and is more difficult to detect. To explore the impact of code correctness, we separately present the detection AUROC for both correct and incorrect codes on the MBPP benchmark when using CodeLlama as Detector LLM in Figure 8.
@@
In Figure 8, we observe that both with our method and other baselines, detecting correct code is more challenging than detecting incorrect code. Nonetheless, our method consistently outperforms other baselines when it comes to correct code.
@@ 6.7. Impact of Code Length
We also investigate how the length of the code snippet impacts the detection performance of the zero-shot synthetic code detectors. We divide the samples in the APPS benchmark into five groups according to the code string length. Each group has an equal number of synthetic and human-written code snippets. The results are plotted in Figure 9.
@@
In general, the performance of all detectors improves as the code length falls within the 0-1000 character range. The $\log p(x)$, LogRank, and Entropy methods achieve optimal detection performance on code snippets between 600-1000 while exhibiting inferior performance on code lengths smaller than 200. DetectGPT’s performance surpasses ours when the code exceeds 1000 characters. However, DetectGPT is inferior to the other baselines for code lengths less than 600, the range where the majority ($70\%$) of code snippets lie. Our method displays strong performance across all ranges and is more robust to changes in code length when compared to the other baselines.
@@ 6.8. Choice of $m$
In our primary experiment, we set the number of rewriting $m$ in the range of $[2,4,8]$ due to limited computational resources. However, increasing $m$ can reduce randomness in code sampling and enhance the accuracy of expectation estimation. To investigate this, we conduct experiments by setting the maximum value of $m$ to 32 on the APPS and MBPP benchmarks (generated by GPT-3.5-turbo) and plotted the detection AUROC against changes of $m$ in Figure 10. The result indicates that the detection performance increases with $m$ with slight fluctuation and saturates around 20 - 32 rewrites.
@@ 7. Conclusion
In this paper, we identified the gap when applying synthetic content detectors designed for general texts to code content, and proposed a novel zero-shot synthetic code detector via code rewriting and similarity measurement. We utilize the similarity between rewritten and original code as an indicator for detecting synthetic code. Our proposed method addresses the unique challenges posed by the massive ”low-entropy” tokens in code domain. Through exhaustive experiments, we uncovered several findings regarding synthetic code detection. The proposed method demonstrates superior performance and robustness in detecting synthetic code compared to other zero-shot baselines.